OpenStudy (anonymous):

(sqrt 2+8) (sqrt 2-8) multiplying radical terms

OpenStudy (anonymous):

\[\sqrt{x}\times \sqrt{y} = \sqrt{xy}\] Does this help?

OpenStudy (anonymous):

ya thnks